# UGG: Unified Generative Grasping

### v1

@@ 3.2 Overview of Our Approach
Given sufficient discriminative guidance, a classifier could be learned to predict whether a candidate’s posture would be able to successfully grasp the target object or not. However, discriminative guidance, e.g., the label of whether a grasp would be successful, is not readily available. To obtain such discriminative labels for training, we resort to the physics simulator, Isaac Gym [43], to judge whether a rendered candidate would compose a successful grasping. This discriminator takes advantage of the diversity of the results generated and significantly improves the success rate of grasping, shown in Sec. 6.
@@ 4.2 Unified Contact Modeling
Figure 3: Subtle object changes can lead to grasping failure, as the corresponding adjustment in latent representation may not capture critical details adequately.
Object and hand encoders are individually trained to facilitate their own shape reconstruction, without aligning together for good grasping. As shown in Fig. 3, a subtle change in the object can cause grasping failure. The corresponding change in latent representation may not be significant enough to capture critical detail. This underscores the need to incorporate contact information into the unified representation.
Previous approaches have taken two main routes to capture this information. Pre-processing methods [34, 15, 22] generate contact information directly before employing optimization algorithms to fit a hand model. The post-processing approach, e.g., GraspTTA [21], involves a two-stage training process to first compute a possible contact map based on an initial hand estimate, and then refine the pose on it. Unfortunately, neither of the existing contact modelings successfully accomplishes the joint generation of both hand parameters and a contact model.
@@ 4.4 Discriminator
Specifically, the evaluation process involves a physical simulator, Isaac Gym[43], to determine the grasping quality. The simulator is not differentiable and, therefore, cannot be implemented as an objective during training. To mitigate this gap between training and evaluation, a lightweight recognition model, termed a Physics Discriminator, is introduced as an efficient component for the purpose of identifying successful samples among the generated postures. This incorporation harnesses the high diversity and propels towards an enhanced success rate of grasping.
A notable challenge arises from the fact that existing datasets inherently contain only valid grasping, without any invalid samples during the discriminator’s training. Fortunately, with the generative nature of the proposed model, adversarial samples with substantial coverage can be obtained from a suboptimal model (e.g., an early-stage model). The ground truths for training are acquired by the physical simulator, with successful grasping as ${\bm{y}}=[0,1]^{T}$, and unsuccessful as ${\bm{y}}=[1,0]^{T}$. Given an object point cloud ${\bm{o}}$ and a generated hand $\hat{{\bm{h}}}$, physical discriminator $D$ is trained with a cross-entropy loss:
@@ 6 Experiment and Evaluation
6.1 Experimental Setup
Dataset. We evaluate our method using the DexGraspNet [71], a challenging large-scale dexterous grasping benchmark. It encompasses 1.32 million dexterous grasps on 5355 objects, following the structure of ShadowHand [57]. Each object is associated with over 200 grasps, and the dataset is curated through an optimization algorithm. Rigorous validation is conducted in the Isaac Gym simulator [43] and on the penetration depth, ensuring the effectiveness and reliability of all grasps. The training set consists of 4229 objects, while the test set consists of 1126 objects among which 236 objects are considered objects from novel categories.
Given the complex nature of this dataset, we employ two subsets to generate results and conduct the ablation study. The first subset, referred to as “20 objects”, comprises a training and testing set featuring 20 carefully selected objects in 12 categories, including two novel objects in the testing set. This subset plays a crucial role in our ablation study and serves as the basis for constructing the object generation model. The second subset, termed “10 bottles”, specifically includes 10 bottles for training the object and the joint generation model. This choice aligns with previous studies on point cloud generation [78, 42], where models are trained per category. Further details on these subsets are provided in the supp. material.
Metrics. In our evaluation, we adhere to the metrics established in DexGraspNet, ensuring a fair comparison with baseline methods. These metrics encompass two key aspects: quality and diversity. To assess grasp quality, three metrics are used: 1) Success rate (%): A grasp is considered successful if it can endure at least one of the six gravity directions in the Isaac Gym simulator and exhibits a maximal penetration depth of less than 0.5 cm. 2) Mean $Q_{1}$: $Q_{1}$ [14] is defined as the radius of the inscribed sphere of ConvexHull ($\cup_{i}w_{i}$), representing the norm of the smallest wrench that can destabilize the grasp. The contact threshold is set to 1cm. 3) Maximal penetration depth (cm): The maximal penetration depth is calculated as the penetration depth from the object point cloud to the hand meshes. Diversity is evaluated through joint angle entropy. The joint motion range is discretized into 100 bins, and all generated samples contribute to estimating a probability distribution. The entropy is then computed as the entropy of these distributions. The reported value represents the mean and standard deviation across all joints.
Baseline Methods. To the best of our knowledge, only two competitive methods have been fully evaluated on DexGraspNet: DDG [33] and GraspTTA [21]. DDG [33] introduces a differentiable $Q_{1}$ metric and supervises a regression-based network to predict grasps end-to-end. The network takes a set of five depth images as input and regresses the pose and joint angles for one grasp. GraspTTA [21] follows a two-stage training strategy. It generates a coarse grasp conditioned on the point cloud of a given object through a CVAE network initially, followed by a contact net to predict a contact map. The contact anchors, along with the penetration energy, are introduced to a test-time adaption stage where the grasp undergoes final refinement.
@@ 6.2 Results of Dexterous Grasping
Quantitative Results. The quantitative results, as depicted in Tab. 1, underscore the superior performance of UGG compared to baseline methods. The performance metrics for the baseline methods are referenced from the DexGraspNet [71] benchmark. UGG achieves the highest success rate (72.7%), $Q_{1}$ (0.063), and the lowest penetration (0.14) among all methods. In terms of diversity, UGG outperforms DDG and GraspTTA significantly, demonstrating superiority in both mean entropy (7.17) and standard deviation (0.07). This highlights the efficacy of the proposed unified diffusion model. In particular, even without the proposed physics discriminator, our method achieves a success rate of 64.1%, only marginally trailing the DDG method. Our approach, drawing insights from both generative and regression-based models, establishes itself as a state-of-the-art performer in terms of both quality and diversity. For seen categories, our method achieves an average success rate of 73.8%, while for novel categories, our model achieves a success rate of 68.1%. These results underscore the strong generalizability of UGG to novel categories.
@@
Qualitative Results. Fig. 4 illustrates the grasps generated by UGG. The top two rows display results for objects within seen categories (though not part of the training set), while the bottom row showcases results for objects falling under novel categories not encountered during training. Each object is accompanied by three distinct generated grasps, highlighting the considerable diversity of the generated grasps. We highlight the depiction of mugs in the middle, emphasizing the model’s capacity to encompass even the handle in its generated grasps. Another noteworthy instance is the headphone in the top right corner, where the generative model exhibits proficiency in devising grasping approaches involving both earpads and the headband.
Additionally, Fig. 5 depicts the generated contact anchors. These anchors signify crucial positions for the grasping task, and our hand model adeptly conforms to these positions. The generated contact anchors demonstrate high diversity, preventing the generated grasps from being confined to fixed poses. Even for unseen objects, such as the bunny car on the right, the contact anchors approach the object and identify valuable grasping points. Some contact anchors may appear close to each other, a natural occurrence based on the relatively strict threshold used in computing the training data. This threshold is set to align with the precise nature of the grasping task.
@@ 6.4 Ablation Study
Physics Discriminator. We conduct an ablation study on the physics discriminator to assess its performance and how the hyper-parameter influences the quantitative results. We consider two settings: 1) Setting different thresholds for predicted probabilities and 2) Selecting a different number of poses by ranking the predicted probabilities. A good discriminator is expected to show an increasing success rate with an increase in the threshold and a decreasing success rate with a larger number of selected poses. We generated a total of $1000$ poses for this study.
@@
Ablation on Hand Encoder and Contact Anchor. To assess the effectiveness of each component in our UGG model, we conduct an ablation study on the generative model. The baseline model comprises a unified generation model without incorporating hand joint encoding or using the contact anchor. Another baseline model includes the hand joint encoder but excludes the use of the contact anchor. Given the complexity of the extensive dataset, we train and test three models on a smaller subset involving 20 objects, with details provided in the appendix. Each model undergoes testing without the physics discriminator to ensure a fair comparison of its generative performance.
Tab. 2 presents results, revealing that the model incorporating both components surpasses the two baseline methods in both success rate and diversity. This superiority is attributed to the effectiveness of the proposed contact anchors, which provide valuable guidance for grasp generation during the generation and optimization phases. The model with the hand encoder does not demonstrate significant difference in success rate but enhances diversity compared to the model without the hand encoder. The encoding of hand joint parameters is akin to a data augmentation method, introducing increased diversity. It should be noted that the baseline method without the hand encoder and contact anchor can be considered as a variant of the method proposed in SceneDiffuser [20] (reproducing SceneDiffuser in this dataset is not feasible). This ablation study highlights the substantial contributions of our proposed encoding schemes and contact modeling to the success of UGG.
Limitation and Failure Case Analysis. Fig. 9 presents a comprehensive exploration of four prominent failure types identified in our generation results. A noteworthy challenge arises when certain generated poses exhibit substantial penetration between the hand and the object (red), contrasting cases where the hand and object become completely detached (green). An intriguing observation within the generation results is the existence of grasps that lean more towards a touch rather than a comprehensive grasp (blue), such results can sometimes even fool our physics discriminator. Additionally, we visualize instances where the discriminator erroneously categorizes generated grasps capable of holding the object as negative samples (yellow). In general, most of the presented cases only require subtle adjustments to be valid. Although our methodology currently stands as the state-of-the-art in dexterous grasping, we acknowledge the need for future endeavors to refine the physics discriminator and strike a more optimal balance between penetration, simulated success, and grasp diversity.
Figure 9: Failure case analysis of UGG. We show four types of failure cases in the generation process. Red: Succeed in the simulator but the penetration is considered too large. Green: Hand detached from the object. Blue: More of a touch instead of a grasp, yet could fool the discriminator. Yellow: A valid grasp from simulation but considered as a failure by the physics discriminator.
@@ 7 Conclusions
This work tackles the task of dexterous grasping. By introducing a novel representation of contact information, Contact Anchor, for the first time, it models objects, hands, and contacts in a unified diffusion process. A physics discriminator is carefully designed to utilize the diversity of a generative model, and further push the success rate of grasping. The combining of all these novelties results in an effective framework that outperforms all state-of-the-art results on both success rate and diversity. In addition, grasping generation from hand to object, or even both from scratch, becomes possible, with the help of the proposed unified design.
Appendix
Appendix A Generative Grasping Model
A.1 Introduction of Diffusion Model
@@
2       ${\bm{ohc}}_{0}=({\bm{o}}_{0},{\bm{h}}_{0},{\bm{c}}_{0})\sim q({\bm{o}}_{0},{\bm{h}}_{0},{\bm{c}}_{0})$ ${\bm{t}}=(t^{o},t^{h},t^{c})\sim\text{Uniform}(\left\{1,\ldots,T\right\})$ ${\bm{\epsilon}}=({\bm{\epsilon}}^{o},{\bm{\epsilon}}^{h},{\bm{\epsilon}}^{c})\sim\mathcal{N}({\bm{0}},\mathbf{I})$ ${\bm{ohc}}_{{\bm{t}}}=\sqrt{\bar{\alpha}_{{\bm{t}}}}{\bm{ohc}}_{0}+\sqrt{1-\bar{\alpha}_{{\bm{t}}}}{\bm{\epsilon}}$ Take gradient step on $\nabla_{\xi}\lVert{\bm{\epsilon}}_{\xi}({\bm{ohc}}_{{\bm{t}}},{\bm{t}})-{\bm{\epsilon}}\rVert^{2}_{2}$
until converged
@@
1 ${\bm{h}}_{T},{\bm{c}}_{T}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ for $t$ = $T,\ldots,1$ do
2       ${\bm{z}}^{h},{\bm{z}}^{c}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ if $t>1$, else ${\bm{z}}^{h},{\bm{z}}^{c}={\bm{0}}$ ${\bm{ohc}}_{t}=({\bm{o}}_{0},{\bm{h}}_{t},{\bm{c}}_{t})$ ${\bm{t}}=(0,t,t)$ ${\bm{\epsilon}}={\bm{\epsilon}}_{\xi}({\bm{ohc}},{\bm{t}})$ ${\bm{h}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{h}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}^{h}\right)+\sigma_{t}{\bm{z}}^{h}$ ${\bm{c}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{c}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}^{c}\right)+\sigma_{t}{\bm{z}}^{c}$
@@ 1 ${\bm{h}}_{T},{\bm{c}}_{T}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ for $t$ = $T,\ldots,1$ do
1 ${\bm{ohc}}_{T}=({\bm{o}}_{T},{\bm{h}}_{T},{\bm{c}}_{T})\sim\mathcal{N}({\bm{0}},\mathbf{I})$ for $t$ = $T,\ldots,1$ do

### v2

@@ Abstract
Keywords:
Dexterous Grasping Contact Representation Generative Model
@@ 3.2 Overview of Our Approach
Given sufficient discriminative guidance, a classifier could be learned to predict whether a candidate’s posture would be able to successfully grasp the target object or not. However, discriminative guidance, e.g., the label of whether a grasp would be successful, is not readily available. To obtain such discriminative labels for training, we resort to the physics simulator, Isaac Gym [46], to judge whether a rendered candidate would compose a successful grasping. This discriminator takes advantage of the diversity of the results generated and significantly improves the success rate of grasping, shown in Sec. 6. Furthermore, we demonstrate that with the bless of high initial diversity and quality, the discriminative information can be utilized to adjust success rate and diversity, a point overlooked in related work based on generative models [48].
@@ 4.2 Unified Contact Modeling
Figure 3: Motivation of a unified contact modeling. (a) Subtle object changes can lead to grasping failure, as the corresponding adjustment in latent representation may not capture critical details adequately. (b) Contact map fails when embedded into a joint generation model. Left: generated noisy contact map (deeper for closer). Right: generated Contact Anchor (yellow) with a grasp and the GT contact map of the generated grasp. Zoom in for better view.
Object and hand encoders are individually trained to facilitate their own shape reconstruction, without aligning together for good grasping. As shown in Fig. 3(a), a subtle change in the object can cause grasping failure. The corresponding change in latent representation may not be significant enough to capture critical detail. This underscores the need to incorporate contact information into the unified representation.
Previous approaches have taken two main routes to capture this information. Pre-processing methods [38, 17, 26] generate contact information directly before employing optimization algorithms to fit a hand model. The post-processing approach, e.g., GraspTTA [24], involves a two-stage training process to first compute a possible contact map based on an initial hand estimate, and then refine the pose on it. Unfortunately, neither of the existing contact modelings successfully accomplishes the joint generation of both hand parameters and a contact model. Meanwhile, our attempts in embedding the contact map modeling into the joint generative model (Fig. 3(b)) yielded pure noise. We attribute it to the unordered nature of the point cloud, which failed to provide a clear pattern for generative contact modeling.
@@ 4.4 Physics Discriminator
Specifically, the evaluation process involves a physical simulator, Isaac Gym[46], to determine the grasping quality. The simulator is not differentiable and, therefore, cannot be implemented as an objective during training. To bridge this gap, we introduce a lightweight recognition model called the Physics Discriminator. This model efficiently identifies successful samples among the generated postures and operates independently from the generative and adaptation modules. We also emphasize its capability of balancing success rate and diversity, an aspect subtlety overlooked by prior methods. This potential becomes evident now when the base model achieves both a commendable success rate and high diversity. Our ablation study in Sec. 6.4 shows that the Physics Discriminator enables control over the generation success rate and diversity.
A challenge arises from existing datasets containing only valid grasps, lacking invalid samples for training the discriminator. Fortunately, our generative model can easily create adversarial samples. Ground truths for training are obtained via the physical simulator, labeling successful grasps as ${\bm{y}}=[0,1]^{T}$ and unsuccessful ones as ${\bm{y}}=[1,0]^{T}$. Given an object point cloud ${\bm{o}}$ and a generated hand $\hat{{\bm{h}}}$, the physical discriminator $D$ is trained with a cross-entropy loss:
@@ 6 Experiment and Evaluation
6.1 Experiment Setup
Dataset. We benchmark our method using the DexGraspNet [76], a challenging large-scale dexterous grasping benchmark. It encompasses 1.32 million dexterous grasps on 5355 objects, following the structure of ShadowHand [61]. Validation is done in the Isaac Gym simulator [46] and on penetration depth, ensuring grasp effectiveness. The training set includes 4229 objects, and the test set has 1126 objects, with 236 from novel categories. Given the complex nature of this dataset, we employ two subsets, namely “20 objects” and “10 bottles”, to generate results and conduct the ablation study, due to the large training time on the entire dataset. We also tested our method on the human-object interaction datasets HO3D [19] and GRAB [69, 5]. Further details on these datasets and results are provided in the appendix.
Metrics. In our evaluation, we adhere to the metrics established in each benchmark, ensuring a fair comparison with baseline methods. These metrics encompass two key aspects: quality and diversity. For all metrics, we follow [76, 24, 25, 38] for implementation. Details of each metric are provided in the supp. material.
Baseline Methods. We compare three methods on the DexGraspNet benchmark dataset: DDG [37], GraspTTA [24], and the generation module in UniDexGrasp [82] (abbr. UDG).
@@ 6.2 Results of Dexterous Grasping
Quantitative Results. The quantitative results, as depicted in Tab. 1, underscore the superior performance of UGG compared to baseline methods.
@@
UGG achieves the highest success rate (72.7%), $Q_{1}$ (0.063), and the lowest penetration (0.14) among all methods. In terms of diversity, UGG outperforms DDG, GraspTTA, and UniDexGrasp significantly, demonstrating superiority in both mean entropy (7.17) and standard deviation (0.07). This highlights the efficacy of the proposed unified diffusion model.
In particular, even without the proposed physics discriminator, our method achieves a success rate of 64.1%, only marginally trailing the DDG method. Our approach, drawing insights from both generative and regression-based models, establishes itself as a state-of-the-art performer in terms of both quality and diversity. For seen categories, our method achieves an average success rate of 73.8%, while for novel categories, our model achieves a success rate of 68.1%. These results underscore the strong generalizability of UGG to novel categories.
@@
Qualitative Results. Fig. 4 illustrates the grasps generated by UGG. The top two rows display results for objects within seen categories (though not part of the training set), while the bottom row showcases results for objects falling under novel categories not encountered during training. Each object is accompanied by three distinct generated grasps, highlighting the considerable diversity of the generated grasps. The depiction of mugs in the middle emphasizes the model’s capacity to encompass even the handle in its generated grasps. Another noteworthy instance is the headphone in the top right corner, where the generative model exhibits proficiency in devising grasping approaches involving both earpads and the headband.
Additionally, Fig. 5 shows the generated contact anchors, which highlight key positions for grasping. Our hand model conforms well to these diverse anchors, avoiding fixed poses. For unseen objects like the bunny car on the right, the anchors effectively identify grasp points. When the contact area between hand and object is small, anchors may appear close due to a strict threshold in training data, which aligns with the precision of the grasping task and does not negatively affect the optimization algorithm.
@@ 6.4 Ablation Study
Figure 9: Four types of failure cases in the generation process of UGG.
Physics Discriminator. We conduct an ablation study on the physics discriminator to assess its performance and how the hyper-parameter helps to balance quality and diversity. We consider two settings: 1) Setting different thresholds for predicted probabilities and 2) Selecting a different number of poses by ranking the predicted probabilities. A good discriminator is expected to show an increasing success rate with an increase in the threshold and a decreasing success rate with a larger number of selected poses. We generated a total of $1000$ poses for this study.
@@
Ablation on Hand Encoder and Contact Anchor. We conducted an ablation study on our UGG model to assess each component’s effectiveness. The baseline model is without hand joint encoding and the contact anchor. Another baseline includes hand joint encoding but not the contact anchor. Due to dataset complexity, we trained and tested three models on a smaller subset of 20 objects (details in the appendix). Each model was tested without the physics discriminator for a fair comparison of generative performance.
@@
Tab. 2 shows that the model with both components outperforms two baselines in success rate and diversity. This superiority is attributed to the effectiveness of the proposed contact anchors, which provide valuable guidance for grasp generation during the generation and optimization phases. The model with the hand encoder improves diversity without significantly affecting the success rate. Encoding hand joint parameters acts like data augmentation, introducing increased diversity. The baseline without the hand encoder and contact anchor is similar to a variant of SceneDiffuser [23]. This ablation study highlights the substantial contributions of our proposed encoding schemes and contact modeling to the success of UGG.
Limitation and Failure Case Analysis. Fig. 9 explores four key failure types in our generation results. One challenge is significant penetration between the hand and object (red), contrasting with cases where they are completely detached (green). We also observe grasps that lean more towards touch than a full grasp (blue), sometimes even fooling our physics discriminator. Additionally, we visualize instances where the discriminator erroneously categorizes successful grasps as failures (yellow). In general, most of the presented cases only require subtle adjustments to be valid. Although our methodology currently stands as the state-of-the-art in dexterous grasping, we acknowledge the need for future endeavors to refine the physics discriminator and strike a more optimal balance between penetration, simulated success, and grasp diversity.
@@ 7 Conclusions
This work tackles the task of dexterous grasping. By introducing a novel representation of contact information, Contact Anchor, for the first time, it models objects, hands, and contacts in a unified diffusion process. A physics discriminator is carefully designed to utilize the diversity of a generative model and further push the success rate of grasping. Combining all these novelties results in an effective framework that outperforms all state-of-the-art results on both success rate and diversity. Additionally, grasping generation from hand to object, or even both from scratch, becomes possible with the proposed unified design.
Acknowledgements
We would like to acknowledge NSF IIS-2047677, HDR-1934932, and CCF-2019844.
Appendix 0.A Generative Grasping Model
0.A.1 Introduction of Diffusion Model
@@
2       ${\bm{ohc}}_{0}=({\bm{o}}_{0},{\bm{h}}_{0},{\bm{c}}_{0})\sim q({\bm{o}}_{0},{\bm{h}}_{0},{\bm{c}}_{0})$
3       ${\bm{t}}=(t^{o},t^{h},t^{c})\sim\text{Uniform}(\left\{1,\ldots,T\right\})$
4       ${\bm{\epsilon}}=({\bm{\epsilon}}^{o},{\bm{\epsilon}}^{h},{\bm{\epsilon}}^{c})\sim\mathcal{N}({\bm{0}},\mathbf{I})$
5       ${\bm{ohc}}_{{\bm{t}}}=\sqrt{\bar{\alpha}_{{\bm{t}}}}{\bm{ohc}}_{0}+\sqrt{1-\bar{\alpha}_{{\bm{t}}}}{\bm{\epsilon}}$
6       Take gradient step on $\nabla_{\xi}\lVert{\bm{\epsilon}}_{\xi}({\bm{ohc}}_{{\bm{t}}},{\bm{t}})-{\bm{\epsilon}}\rVert^{2}_{2}$
7
8until converged
@@ 3       ${\bm{t}}=(t^{o},t^{h},t^{c})\sim\text{Uniform}(\left\{1,\ldots,T\right\})$
1 ${\bm{h}}_{T},{\bm{c}}_{T}\sim\mathcal{N}({\bm{0}},\mathbf{I})$
2 for $t$ = $T,\ldots,1$ do
3       ${\bm{z}}^{h},{\bm{z}}^{c}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ if $t>1$, else ${\bm{z}}^{h},{\bm{z}}^{c}={\bm{0}}$
4       ${\bm{ohc}}_{t}=({\bm{o}}_{0},{\bm{h}}_{t},{\bm{c}}_{t})$
5       ${\bm{t}}=(0,t,t)$
6       ${\bm{\epsilon}}={\bm{\epsilon}}_{\xi}({\bm{ohc}},{\bm{t}})$
7       ${\bm{h}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{h}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}^{h}\right)+\sigma_{t}{\bm{z}}^{h}$
8       ${\bm{c}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{c}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}^{c}\right)+\sigma_{t}{\bm{z}}^{c}$
9
@@ 6       ${\bm{\epsilon}}={\bm{\epsilon}}_{\xi}({\bm{ohc}},{\bm{t}})$
1 ${\bm{ohc}}_{T}=({\bm{o}}_{T},{\bm{h}}_{T},{\bm{c}}_{T})\sim\mathcal{N}({\bm{0}},\mathbf{I})$
2 for $t$ = $T,\ldots,1$ do
3       ${\bm{z}}=({\bm{z}}^{o},{\bm{z}}^{h},{\bm{z}}^{c})\sim\mathcal{N}({\bm{0}},\mathbf{I})$ if $t>1$, else ${\bm{z}}={\bm{0}}$